# Computing almost commuting bases of ODOs and Gelfand-Dikey hierarchies

### v1

@@ 5.2 Fith order operators and their hierarchy
For each $m\geq 1$, it holds
$[L_{5},P_{m}]=H_{m,3}(U)\partial^{3}+H_{m,2}(U)\partial^{2}+H_{m,1}(U)\partial+H_{m,0}(U),$
@@
The system of the Gelfand-Dikii hierarchy at level $m=9$ in (27) is ,
@@
We computed $H_{j,5-i}$ for $1\leq j\leq 9$ and $i=2,\ldots,5$. The results of these computations can be found in the repository da_wilson, where we have collected several computations of these hierarchies. In particular, we refer to the folder latex, where the LaTeX representation of these polynomials $H_{i,j}(U)$ can be found in the files with name (5_i)[H_j].tex. A further description of this repository can be found in Section 6.
@@ 6 Implementation
In this paper we have revisit the theory of almost commuting differential operators and the definition of the homogeneous basis for these operators. In order to allow future computations, we have implemented Algorithm 1 in the computer algebra system SageMath [19], in the package dalgebra, a set of tools implemented in SageMath by the first author dedicated to Differential Algebra. We present in this section the main features of this implementation, showing also the time spent for these computations and presenting a public repository where the results can be retrieved in several formats.
@@ 6.1 Implementation of Algorithm 1
  • The list of $(n-1)$ differential polynomials $H_{m,i}(U)\in\mathbb{Q}\{U\}$ such that
    $[L_{n},P_{m}]=H_{m,0}(U)+H_{m,1}(U)\partial+\ldots+H_{m,n-2}(U)\partial^{n-2}.$
@@
It can be appreciated that the times are still reasonable (not bigger than 120 seconds). We would like to remark that the bigger result (in space) is the case with $n=7$ and $m=13$, where all data structures takes 4MB of disk space. Mathematically speaking, this result ($n=7$, $m=13$) contains the polynomial $P_{13}$ for $L_{7}$ with 830 monomials, and the 6 differential polynomials $H_{13,i}$ all with degree 7 and ranging from 744 up to 5279 monomials. The script that generates the figure (i.e., measures the times and generates the image) can be found in folder experiments in the aforementioned repository, in the file almost_commuting.sage.
@@ 6.2 Results available
We refer to the README in this repository for explanation on advice on how to load data from these files into the different systems. Also, the script for generating all the Maple and TeX files is included in this repository.
Acknowledgements. The authors are partially supported by the grant PID2021-124473NB-I00, “Algorithmic Differential Algebra and Integrability” (ADAI) from the Spanish MICINN. The first author was partially supported by the Poul Due Jensen Grant 883901.

### v2

@@ 5.2 Fifth order operators and their GD hierarchy
For each $m\geq 1$, it holds that
$P_{m},L_{5}]=H_{m,3}(U)\partial^{3}+H_{m,2}(U)\partial^{2}+H_{m,1}(U)\partial+H_{m,0}(U),$
@@
The system of the Gelfand-Dickey hierarchy at level $m=9$ in (28) is ,
@@
We computed $H_{j,5-i}$ for $1\leq j\leq 9$ and $i=2,\ldots,5$. The results of these computations can be found in the repository da_wilson, where we have collected several computations of these hierarchies to provide a data set that can be used by researchers in Mathematics and Physics. In particular, we refer to the folder latex, where the LaTeX representation of these polynomials $H_{i,j}(U)$ can be found in the files with name (5_i)[H_j].tex. A further description of this repository can be found in Section 6.
@@ 6 Implementation
In this paper, we revisit the theory of almost commuting differential operators and the definition of the homogeneous basis for these operators. In order to allow future computations, we have implemented Algorithm 1 in the computer algebra system SageMath [17], in the package dalgebra, a set of tools implemented in SageMath by the first author and dedicated to Differential Algebra. We present in this section the main features of this implementation, showing the time spent for these computations and presenting a public repository where the results can be retrieved in several formats.
Although a complexity analysis of Algorithm  1 was not performed, we would like to point out that the complexity would depend on: the multiplication of differential operators in Step 2; and the complexity of the integration algorithm used in Step 4, the instruction $\texttt{INTEG}(e_{m,i})$ of Algorithm 2. It is important to emphasize that the system $\mathbb{E}_{m,n}$ of Step 3 is already triangular, as proved in Theorem 4.6, so there is no triangulation algorithm involved.
@@ 6.1 Implementation of Algorithm 1
  • The list of (n1)(n-1) differential polynomials $H_{m,i}(U)\in\mathbb{Q}\{U\}$ such that
    $P_{m},L_{n}]=H_{m,0}(U)+H_{m,1}(U)\partial+\ldots+H_{m,n-2}(U)\partial^{n-2}.$
@@
It can be appreciated that the times are still reasonable (not bigger than 120 seconds). We would like to remark that the biggest result (in space) is the case with $n=7$ and $m=13$, where all data structures take 4MB of disk space. Mathematically speaking, this result ($n=7$, $m=13$) contains the polynomial $P_{13}$ for $L_{7}$ with 830 monomials, and the 6 differential polynomials $H_{13,i}$ all with degree 7 and ranging from 744 up to 5279 monomials. The script that generates the figure (i.e., measures the times and generates the image) can be found in folder experiments in the aforementioned repository, in the file almost_commuting.sage.
@@ 6.2 Results available
We refer to the README in this repository for advice on how to load data from these files into the different systems. Also, the script for generating all the Maple and TeX files is included in this repository.
[…]
Acknowledgements All authors are partially supported by the grant PID2021-124473NB-I00, “Algorithmic Differential Algebra and Integrability” (ADAI) from the Spanish MICINN. R.D. is partially supported by the Poul Due Jensen Grant 883901.